OpenStudy (anonymous):

. The brakes of a car are applied to give it an acceleration of -3 m/s2. The car comes to a stop in 5 s. What was its speed when the brakes were applied?

OpenStudy (anonymous):

Vf =Vi +at Vf=0 (stopped) -Vi = -3(5)